# Pushing LLMs to Their Logical Reasoning Bound: The Role of Data Reasoning Intensity

### v3

@@ Abstract
Recent advances in large language models (LLMs) highlight the importance of training data structure and quality in shaping reasoning behavior. However, most existing approaches focus on transforming data formats while neglecting the internal reasoning complexity of training samples, leaving the reasoning potential of data underexplored and underutilized. In this work, we posit that LLM logical reasoning performance is jointly constrained by the potential of the training data and the cognitive capacity of the model. To make this relationship measurable, we introduce Data Reasoning Intensity (DRI), a novel metric that quantifies the latent logical reasoning complexity of samples by decomposing and aggregating their logical structures. This allows us to analyze how well current LLMs utilize logical reasoning signals and identify performance gaps relative to data potential. Based on this insight, we introduce a re-cognizing optimization strategy that systematically enhances the logical reasoning intensity of training data. Rather than increasing data volume, our method re-optimizes existing samples to better align with the LLM’s logical reasoning boundary. Extensive experiments show that our approach significantly improves performance and generalization over data-centric strategies. We further validate our method under a reinforcement learning framework. Our results indicate that prioritizing reasoning complexity in data rather than sheer scale or superficial form is essential to realizing LLMs’ full cognitive potential.
@@ 1 Introduction
The reasoning ability of LLMs (OpenAI, 2023; DeepSeek-AI, 2024; Cheng et al., 2025; Chen et al., 2025d; Gao et al., 2025b; Ke et al., 2025; Zhou et al., 2024; Qiao et al., 2023) has emerged as a core metric for evaluating their cognitive alignment with human-like problem-solving. With the breakthrough of LLMs in logical reasoning tasks (Chen et al., 2025c; Li et al., 2025b; Feng et al., 2025; Wang et al., 2025a), optimizing the cognitive abilities of models from the perspective of training data has become the mainstream paradigm (Wu et al., 2025; Kandpal & Raffel, 2025; Peng et al., 2025; Prystawski et al., 2023; Chen et al., 2024; Kim et al., 2025). By reconstructing the cognitive expression form of the training data rather than simply expanding the data scale, it has demonstrated the crucial influence of data quality on the boundary of LLM’s logical capabilities.
Figure 1: The illustrative visualization of data potential and LLM logical reasoning performance. The horizontal axis denotes data potential and the vertical axis denotes an LLM’s logical-reasoning performance. Each curve corresponds to a specific LLM (e.g., different sizes) performance bound under a fixed architecture. As the potential of data increases, the performance of LLMs usually improves, but eventually it will reach an upper limit determined jointly by the model’s capacity and the limitations of the data.
In this paper, we aim to explore LLM’s logical reasoning ability from the data-centric perspective. Concretely, advanced methods such as DeepSeek-AI et al. (2025), indicate that it is not more data, but more complex and logically structured data that can better stimulate the reasoning potential of LLMs. Zhou et al. (2023); Ye et al. (2025) demonstrate that the structural guidance of training data is more effective than the volume of data in shaping the capabilities and behavioral patterns of LLMs. Indeed, the complexity of data can effectively stimulate and enhance the reasoning ability of LLMs only when the data has a reasonable structure, is close to the model’s experience, and is near the boundary of its reasoning (Chen et al., 2024; Bai et al., 2025; Bi et al., 2024; Prystawski et al., 2023; Liu et al., 2025). Similarly, Peng et al. (2025) shows that the quality structure of training data plays a decisive role in reasoning ability, and precise data selection is better than simple data increment.
However, although existing research has made some progress, the current paradigm has fundamental bottlenecks: its optimization only stays at the superficial level of transforming the form of data expression, while neglecting the in-depth exploration of the data’s internal reasoning complexity. Specifically, existing methods face two major challenges:
  • Lack of efficient quantification: Existing methods lack accurate measures of the logical reasoning complexity contained in data. For LLM’s logical reasoning, logical nodes are crucial components, as they are closely related to the depth, breadth, and accuracy of the reasoning process. As a result, estimating the inherent reasoning complexity of the data itself becomes extremely challenging.
  • Blind analysis and optimization: Although some complexity metrics do exist, few studies have analyzed the impact of training data from the perspective of the logical reasoning bounds of LLM. At the same time, how to leverage these findings in a reverse manner to guide training optimization and proactively enhance LLM’s reasoning abilities remains a key challenge.
To make this gap measurable, we introduce Data Reasoning Intensity (DRI), a practical metric that disassembles the reasoning pipeline into logically coherent steps and re-aggregates them into a single scalar that captures the complexity of the entire logical reasoning chain. Additionally, we suppose that the logical reasoning performance of LLMs is constrained by two bounds: the potential of the training data and the cognitive potential of the models themselves. As illustrated in Figure 1, current LLMs remain far from an ideal state in utilizing available reasoning data—indicating that they have not yet reached the ceiling of either data potential or cognitive capacity. By applying our proposed metric to the training corpora of logical‐reasoning benchmarks, we observe measurable accuracy gains within the same dataset, confirming that the attainable performance ceiling has not been reached. Guided by these observations, we introduce a re-cognizing optimization strategy that systematically enhances the reasoning signals in training data, thereby pushing LLMs closer to their potential logical reasoning performance.
Our main contributions are as follows:
  • We propose a novel metric for evaluating data reasoning potential and analyze the logical reasoning capabilities of LLMs from a data-centric perspective. Our analysis reveals that the training data still holds untapped potential and that current LLMs are far from reaching their performance ceiling.
  • We propose a novel optimization approach, the re-cognizing optimization strategy, which reshapes and enhances the logical reasoning abilities of LLMs across diverse training samples, thereby pushing them closer to their reasoning bound.
  • We validate the existence and rationality of our quantitative metric on logical reasoning tasks. Furthermore, we explore the effectiveness of our proposed optimization strategy, demonstrating its superiority over existing data-centric training paradigms. We also validate its efficacy under a reinforcement learning framework.
@@ 2.1 Data-Centric Reasoning for LLMs
Recent work increasingly emphasizes that, beyond model architecture and algorithm design, the structure, provenance, and difficulty of training data play a critical role in shaping LLM reasoning performance (Chen et al., 2025c; Qu et al., 2025; Sui et al., 2025; Ruis et al., 2025). This has led to a data-centric shift in LLM development, where efforts focus on improving data composition and quality to enhance reasoning ability (Wu et al., 2025; Jin et al., 2024; Wang et al., 2025b). Many studies highlight that higher-quality, logically structured examples—rather than simply more data—yield better generalization and reasoning performance (Peng et al., 2025; Yu et al., 2025; Li et al., 2025a; Wettig et al., 2024; Zhao et al., 2024; Yu et al., 2024). In particular, aligning sample difficulty with model capability is identified as key to effective training (Gao et al., 2025a). Instruction tuning research supports this view, showing that both prompt quality and exposure timing influence reasoning emergence (Qingsong et al., 2025; Kim & Lee, 2024). Meanwhile, Kandpal & Raffel (2025) highlight the often-overlooked human labor cost in curating such training data.
@@ 2.2 Reasoning Emergence for LLMs
Complementary to these data-centric strategies, other studies investigate how reasoning capabilities emerge from the interaction between model cognition and structured input. Stepwise reasoning, for example, has been proposed as an emergent property of sequential data exposure rather than a fixed architectural feature (Prystawski et al., 2023). Furthermore, only data within a suitable complexity range appears to effectively stimulate reasoning behavior (Bi et al., 2024), suggesting that model performance is bounded by cognitive processing capacity. A number of empirical studies further show that small, carefully curated datasets often outperform larger but noisier corpora in supporting reasoning skills (Ma et al., 2025; Ye et al., 2025; Morishita et al., 2024; Wang et al., 2025c; Yang et al., 2025; Hua et al., 2025), reinforcing the value of reasoning supervision that is both selective and structurally rich. Prior work underscores data’s role in LLM reasoning but often lacks precise difficulty metrics and clear goals. We introduce DRI, a unified score for reasoning potential,and re-cognizing optimization, which emphasizes high DRI examples while preserving diversity to enhance LLM’s logical reasoning performance.
3 Methodology
Inspired by the Roofline Model (Williams et al., 2009; Cao et al., 2025), we frame LLM logical reasoning performance as an efficiency ratio between data-driven reasoning potential and model-intrinsic cognitive cost. For a model $\mathcal{M}$ evaluated on a dataset $\mathcal{D}$, the effective reasoning capability $\eta$ is defined as
$\eta(\mathcal{M},\mathcal{D})=\frac{E(\mathcal{D})}{C(\mathcal{M})}$ (1)
where $E(\mathcal{D})$ quantifies the latent logical reasoning demand encoded in the dataset—such as compositional complexity, multi-step inference depth, or symbolic abstraction—while $C(\mathcal{M})$ represents the model’s intrinsic cognitive cost, encompassing architectural capacity, parameter scale, and reasoning FLOPs. Because $C(\mathcal{M})$ reflects hardware-bound and architecture-bound properties that are often costly or slow to modify, the most practical path to improving $\eta$ lies in enriching $E(\mathcal{D})$ through carefully designed, reasoning-intensive data.
Accordingly, we introduce the DRI score to quantify and maximize $E(\mathcal{D})$. It operates in two stages: first, we decompose each sample’s structured reasoning trace into its logical components; second, we compute a DRI score that quantifies the reasoning potential embedded in that structured chain. By focusing on quantifying and raising $E(\mathcal{D})$, our approach unlocks latent training-data potential and drives enhanced LLM reasoning performance. Detailed definitions, explanations and proofs are provided in the Appendix D.
Figure 2: The overall framework. Top: We first extract logical elements from each example and perform combinatorial reasoning to derive a Data Reasoning Intensity (DRI) score. Bottom: We then analyze the performance distribution across DRI levels on multiple datasets. Based on this, we propose the re-cognizing optimization strategy: the first stage reshapes the model’s recognition of reasoning patterns, and the second enhances its logical reasoning capability, thereby improving overall performance.
3.1 Decomposition and Combinatorial Reasoning
LLM logical reasoning begins as a collection of logical elements (Chen et al., 2025c), so we hypothesize that the reasoning process of LLMs involves structured reasoning. We first deconstruct the logical structure of the data and then conduct structural reasoning based on the deconstructed logical elements. The process begins by decomposing each question $Q$ into fundamental logical components via a distillation function $f$, implemented using LLMs such as DeepSeek-V3 or GPT-4o. We then extract the three core logical elements:
$f(Q)\Rightarrow\{\mathcal{P},\mathcal{C},\mathcal{E}\}$ (2)
where $\mathcal{C}$ is the constant set (e.g., "YoungBoy") and $\mathcal{P}$ is the predicates set capturing relationships (e.g., IsYoung(x)). $\mathcal{E}$ is the logic structure sets combining predicates and constants (such as first-order logic expressions).
Based on logical deconstruction, we need to perform combinatorial reasoning on the obtained logical elements. we define a combinatorial reasoning function $F$, implemented with an LLM, that leverages the distilled logical elements $\{\mathcal{P},\mathcal{C},\mathcal{E}\}$ and the candidate answer set $A$ to produce the precondition structure $\bar{\mathcal{E}}$ and the reasoning step sequence $\mathcal{S}$:
$\text{Trajectory}={F}(Element,A)\Rightarrow\{\bar{\mathcal{E}},\mathcal{S}\}$ (3)
where $\bar{\mathcal{E}}$ is the precondition logic structure and $\mathcal{S}$ is the single reasoning chain required for the reasoning trajectory. Detailedly, each reasoning node $s_{k}\in S$ contains:
$s_{k}=(\#\text{Operations}_{k},D_{\text{nest}}^{k},\text{Expression}_{k})$ (4)
where $\text{Operations}_{k}$ captures the logical operators (AND/OR/NOT), $D_{\text{nest}}^{k}$ is the expression’s nesting depth, and $\text{Expression}_{k}$ is its formal representation.
3.2 Data Reasoning Intensity
Logical Intensity Quantification
Following logical decomposition, we compute a context score $S_{\text{ctx}}$ from three dimensions extracted from the question’s logical elements:
$S_{\text{ctx}}=\underbrace{|\mathcal{E}_{i}|\cdot\bar{D}^{2}}_{\text{Structural}}+\underbrace{|\mathcal{P}_{i}|}_{\text{Predicate}}+\underbrace{|\mathcal{C}_{i}|}_{\text{Constant}}$ (5)
where $|\mathcal{E}_{i}|$ counts logical expressions, $\bar{D}$ is their average nesting depth (calculated via parse tree analysis), $|\mathcal{P}_{i}|$ and $|\mathcal{C}_{i}|$ tally unique predicates and constants respectively.
Reasoning Intensity Quantification
Building on combinatorial reasoning, each answer option’s score $S_{\text{opt}}^{(l)}$ combines its precondition intensity with step‐by‐step deduction intensity:
$S_{\text{opt}}^{(l)}\;=\;\underbrace{|\mathcal{R}_{l}|\cdot\bar{D}_{l}^{2}}_{\text{Preconditions}}\;+\;\sum_{k=1}^{T_{l}}\underbrace{\bigl(1+\#\text{Operations}_{l,k}\bigr)\;D_{l,k}^{2}}_{\text{Step }k}$ (6)
where $l\in\{1,\dots,L\}$ indexes the $L$ answer options, $\mathcal{R}_{l}$ is the set of precondition expressions for option $l$ with average nesting depth $\bar{D}_{l}$, and $T_{l}$ is the number of reasoning steps for option $l$. Each step $k$ contributes according to its operator count $\text{Operations}_{l,k}$ and nesting depth $D_{l,k}$.
Since decomposition and combinatorial reasoning metrics capture different facets of DRI and live on disparate scales, we first merge them into a raw intensity measure $S_{\text{raw}}$:
$S_{\text{raw}}$ $=S_{\text{ctx}}+\sum_{l=1}^{L}S_{\text{opt}}^{(l)}$ (7)
To bound the resulting range within [0,1] and obtain the final DRI score $S$, we apply logarithmic compression followed by sigmoid normalization:
$S$ $=\sigma\left(\gamma\cdot\frac{\log(S_{\text{raw}}+1)-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta\right)$ (8)
where $\sigma$ denotes the sigmoid function, $\mu/\sigma^{2}$ are dataset statistics, and parameters ($\gamma=1,\beta=0,\epsilon=10^{-5}$) ensure stable [0,1] normalization.
Figure 3: Effectiveness verification for DRI. Sample counts (bars, left axis) and model error rates (lines, right axis) are shown across DRI score bins. All three panels share the same layout: the x‐axis divides the score range into uniform intervals, the bar height indicates the number of examples per interval, and the overlaid line traces the error rate. (a) Training‐set distribution and error. (b) Original test‐set distribution. (c) Balanced test‐set distribution.
4 Experimental Settings
In this section, we conduct experiments to address the following research questions:
@@ 4 Experimental Settings
    RQ1: Is the DRI score effective in enhancing the logical reasoning performance of LLMs?
@@
    RQ2: How can DRI be effectively leveraged to improve LLMs’ performance during training?
@@
    RQ3: Is the re-cognizing optimization method effective in boosting logical reasoning performance?

### v4

# Logical Structure as Knowledge: Enhancing LLM Reasoning via Structured Logical Knowledge Density Estimation

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) are increasingly attributed to training data quality rather than mere parameter scaling. However, existing data-centric paradigms often equate quality with factuality or diversity and ignore the internal logical complexity of training samples. In this work, we propose that natural language harbors Structured Logical Knowledge manifested through entailment relationships and logical topologies. To quantify this, we introduce Structured Logical Knowledge Density (SLKD), a novel metric that measures logical information content by decomposing natural language into executable predicates and logical primitives. Our analysis reveals a significant logical disparity in current datasets where sparse logical signals predominate. Consequently, we propose a density-aware re-cognizing optimization strategy that prioritizes high-density logical samples to enhance with the LLM’s reasoning ability. Extensive experiments demonstrate that our approach enhances reasoning performance and generalization without increasing total data volume. These results, further validated within a reinforcement learning framework, suggest that elevating logical density is more critical than expanding data scale for realizing the full cognitive potential of LLMs. The released code is available in the Appendix C.
Logical Structure as Knowledge: Enhancing LLM Reasoning via Structured Logical Knowledge Density Estimation
Zhen Bi1,2,3*, Zhenlin Hu1,2thanks: EqualContribution., Xueshu Chen1,2, Mingyang Chen4, Cheng Deng3,5, Yida Xue6, Zhen Wang1,2, Qing Shen 1,2, Ningyu Zhang6, Jungang Lou1,2thanks: CorrespondingAuthor., 1Huzhou University, 2Zhejiang Key Laboratory of Intelligent Education Technology and Application, 3Banbu AI Foundation, 4Baichuan Inc, 5University of Edinburgh, 6Zhejiang University bizhen_zju@zju.edu.cn
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) are attributed not merely to the scaling of parameters but significantly to the enhancement of training data quality. Within the data-centric paradigm (Brown et al., 2020; Raffel et al., 2020; Hoffmann et al., 2022; Zhou et al., 2023), a consensus has emerged: high-quality data serves as the cornerstone of LLM reasoning capabilities.
Figure 1: Comparison between the current methods and our method. The text on the left contains entities and relationships, but it is logically quite straightforward. The logical structure of the text on the right is very dense, but it contains many meaningless symbols.
However, the precise definition of data quality remains elusive. Prevailing research predominantly focuses on dimensions such as factuality and diversity(Hua et al., 2025; Wang et al., 2025a). From this perspective, knowledge is often equated with entities and facts stored within the text, and we argue that this definition is incomplete. Beyond declarative facts, natural language implicitly harbors another crucial form of knowledge: Structured Logical Knowledge. This knowledge manifests not as isolated entities, but as entailment relationships between predicates, conditional constraints, and logical topological structures.
Although there are currently some studies on the complexity of training data for LLMs (Prystawski et al., 2023; Bi et al., 2024), we have long lacked an effective methodology to quantify the precise knowledge density contained within data. Existing analytical approaches typically assess knowledge richness by analyzing factual triples implicitly contained in the text or the underlying syntactic trees. However, they struggle to model the distinctions in knowledge density inherent to texts featuring complex deductive reasoning (shown in Figure 1). This absence of appropriate metrics often leads to an indiscriminate expansion of data volume, which may overlook the critical role of logical structure density as a decisive variable in the emergence of model reasoning capabilities.
To bridge this theoretical gap, we propose regarding logical structure as a measurable form of knowledge. We introduce a novel metric: Structured Logical Knowledge Density (SLKD). Distinct from traditional metrics for text knowledge complexity, SLKD does not focus on the factual knowledge contained in the data, but instead delves into the logical core. By decomposing natural language into executable predicates and logical primitives, and calculating the topological complexity of their combinatorial reasoning processes, it objectively measures the logical information content inherent in a sample.
Leveraging SLKD, we identify a significant logical disparity in existing training data distributions—a vast amount of data provides only sparse logical knowledge, while the importance of high-SLKD samples is overlooked. To address this, we propose a density-aware re-cognizing optimization strategy. Instead of merely pursuing data scale, this strategy reshapes the training process according to SLKD: it first establishes a foundational logical schema via extensive density-based data, and subsequently concentrates computational resources on high-density samples, thereby enabling the model to efficiently acquire complex logical entailment patterns. The main contributions are as follows:
  • We propose the perspective of Logical Structure as Knowledge, highlighting the oversight of logical knowledge density in current data estimation paradigms.
  • We introduce SLKD, an objective metric for quantifying the complexity of internal logical entailment, capable of precisely identifying "high logical value regions" within data.
  • We design an optimization strategy based on SLKD that significantly enhances LLM performance on logical reasoning tasks without increasing the total data volume. This confirms that the key to improving model reasoning lies in elevating the logical knowledge density of training data, rather than merely expanding its scale.
@@ 2.1 Data-Centric Reasoning for LLMs
Recent work increasingly emphasizes that, beyond model architecture and algorithm design, the structure, provenance, and difficulty of training data critically shape LLM reasoning performance (Chen et al., 2025c; Qu et al., 2025; Sui et al., 2025; Ruis et al., 2025). This data-centric shift prioritizes data composition and quality over sheer volume to enhance reasoning ability (Wu et al., 2025; Jin et al., 2024; Wang et al., 2025b). Many studies highlight that higher-quality, logically structured examples—rather than simply more data—yield better generalization and reasoning performance (Peng et al., 2025; Yu et al., 2025; Li et al., 2025; Wettig et al., 2024; Zhao et al., 2024; Yu et al., 2024). In particular, aligning sample difficulty with model capability is identified as key to effective training (Gao et al., 2025). Instruction tuning research supports this view, showing that both prompt quality and exposure timing influence reasoning emergence (Qingsong et al., 2025; Kim and Lee, 2024). Meanwhile, Kandpal and Raffel (2025) highlight the often-overlooked human labor cost in curating such training data.
Figure 2: The overall framework. Top: Logical structure as knowledge. We first extract logical elements from each example and perform combinatorial reasoning to derive Structured Logical Knowledge Density (SLKD) score. Bottom: SLKD-based optimization method. We then analyze the performance distribution across SLKD levels on multiple datasets. Based on this, we propose the re-cognizing optimization strategy: the first stage reshapes the model’s recognition of reasoning patterns, and the second enhances its logical reasoning capability, thereby improving overall performance.
@@ 2.2 Reasoning Emergence for LLMs
Complementing these data-centric strategies, other studies explore how reasoning emerges from the interaction between cognition and structured input. For instance, stepwise reasoning is viewed as an emergent property of sequential data exposure rather than a fixed architectural feature (Prystawski et al., 2023). Furthermore, only data within a suitable complexity range appears to effectively stimulate reasoning behavior (Bi et al., 2024), suggesting that model performance is bounded by cognitive processing capacity. Empirical evidence further shows that carefully curated datasets often outperform larger but noisier corpora in supporting reasoning skills (Ma et al., 2025; Ye et al., 2025; Morishita et al., 2024; Wang et al., 2025c; Yang et al., 2025; Hua et al., 2025), reinforcing the value of reasoning supervision that is both selective and structurally rich. Prior work underscores data’s role in LLM reasoning but often lacks precise difficulty metrics and clear goals. We introduce SLKD, a unified score for reasoning potential, and re-cognizing optimization, which emphasizes high SLKD examples while preserving diversity to enhance LLM’s logical reasoning performance.
3 Logical Structure as Knowledge
In this section, we introduce Structured Logical Knowledge Density (SLKD). The calculation of SLKD involves transforming unstructured natural language into structured predicate logic representations and then evaluating the topological complexity of the resulting reasoning chains.
3.1 Structure Decomposition
The first step in quantifying SLKD is to decouple the logical form from the natural language surface. We treat the reasoning process as a manipulation of executable predicates. Given a data sample $Q$, we employ a distillation function $f$ (implemented via advanced LLMs) to decompose the text into three fundamental logical primitives:
$f(Q)\Rightarrow\{P,C,E\}$ (1)
@@ 3.1 Structure Decomposition
    Constants: $C$ is the set of entities or values involved in the reasoning context (e.g., specific objects, names).
@@
    Predicates: $P$ is the set of relations or properties binding the constants (e.g., $\texttt{IsParentOf}(x,y)$, $\texttt{Greater}(a,b)$).
@@
    Logical Expressions: $E$ is the set of explicit logical constraints or rules from the context, formalized as first-order logic expressions.
This decomposition converts the fuzzy information of natural language into a rigorous predicate knowledge base, serving as the substrate for subsequent density analysis.
3.2 Combinatorial Topology Analysis
The density of logical knowledge arises from how these primitives combine to form valid entailment paths, we define the combinatorial topology analysis process to reconstruct the logical trajectory required to derive the answer from the primitives. Using a function $F$, we generate the reasoning topology, which consists of the precondition structure $\bar{E}$ and a sequence of derivation steps $S$:
$\text{Topology}=F\left(\text{Primitives},A\right)\Rightarrow\{\bar{E},\text{S}\}$ (2)
where $A$ is answer set, $\bar{E}$ is the precondition logic structure and $S$ is the single reasoning chain required for the reasoning trajectory. Detailedly, each reasoning node $s_{k}\in S$ contains:
$s_{k}=\left(\#\text{Operations}_{k},D^{\text{nest}}_{k},\text{Expression}_{k}\right)$ (3)
where $\text{Operations}_{k}$ counts the logical operators (AND/OR/NOT), $D^{\text{nest}}_{k}$ represents the nesting depth of the logical expression, and Expression_k is its formal representation.
The nesting depth is a critical parameter, as it reflects the hierarchical complexity of the entailment relationship.
Figure 3: Effectiveness verification for SLKD. Sample counts (bars, left axis) and model error rates (lines, right axis) are shown across SLKD score bins. All three panels share the same layout: the x-axis divides the score range into uniform intervals, the bar height indicates the number of examples per interval, and the overlaid line traces the error rate. (a) Training-set distribution and error. (b) Original test-set distribution. (c) Balanced test-set distribution.
3.3 Quantifying Structured Logical Knowledge Density (SLKD)
We quantify the density of logical knowledge by aggregating the topological complexity of the context and the derivation path. This measurement encompasses contextual logical density for premise structures, inferential topology density for inferential depth, and SLKD score aggregation for synthesizing these dimensions into a unified metric.
3.3.1 Contextual Logical Density
First, we measure the static logical information contained in the problem description. The Context Density $S_{\text{ctx}}$ is calculated as:
$S_{\mathrm{ctx}}=\underbrace{|{E}_{i}|\cdot\bar{D}^{2}}_{\text{Structural}}+\underbrace{|{P}_{i}|}_{\mathrm{Predicate}}+\underbrace{|{C}_{i}|}_{\mathrm{Constant}}$ (4)
where $|E_{i}|$ counts logical expressions, $\bar{D}$ is their average nesting depth (calculated via parse tree analysis), $|P_{i}|$ and $|C_{i}|$ tally unique predicates and constants respectively.
Here, the term $\bar{D}^{2}$ applies a quadratic penalty to the nesting depth, reflecting the non-linear increase in cognitive load required to encode deeply nested logical structures. This formulation ensures that complex conditional dependencies contribute significantly more to the density score than simple linear facts.
3.3.2 Inferential Topology Density
Next, we quantify the dynamic complexity of the entailment path required to validate a specific option. For an option $l$, its inferential topology density $S^{(l)}_{\text{opt}}$ is defined as:
$S_{\mathrm{opt}}^{(l)}=\underbrace{|{R}_{l}|\cdot\bar{D}_{l}^{2}}_{\text{Preconditions}}+\sum_{k=1}^{T_{l}}\underbrace{\left(1+\#\text{Operations}_{l,k}\right)D_{l,k}^{2}}_{\mathrm{Step~}k}$ (5)
where $l\in\{1,...,\text{L}\}$ indexes the L answer options, $R_{l}$ is the set of precondition expressions for option $l$ with average nesting depth $\bar{D_{l}}$, and $T_{l}$ is the number of reasoning steps for option l. Each step $k$ contributes according to its operator count $\text{Operations}_{l,k}$, and nesting depth $D_{l,k}$.
This component aggregates the structural weight of each derivation step. By summing over the sequence $T_{l}$, we capture the total "logical work" encoded in the reasoning chain.
3.3.3 SLKD Score Aggregation
The final SLKD $S$ is obtained by combining the contextual and derivational densities. To facilitate its use as a weighting factor in training, we normalize the raw score into a probabilistic range $[0,1]$:
$S_{\text{raw}}=S_{\text{ctx}}+\sum_{l=1}^{L}S^{(l)}_{\text{opt}}$ (6)
$S=\sigma\left(\gamma\cdot\frac{\log(S_{\text{raw}}+1)-\mu}{\sqrt{\delta^{2}+\epsilon}}+\beta\right)$ (7)
where $\sigma$ denotes the sigmoid function, $\mu$ and $\delta$ are dataset statistics, and parameters ($\gamma=1$$\beta=0$$\epsilon=10^{-5}$) ensure stable $[0,1]$ normalization.
This normalized SLKD score serves as a proxy for the richness of reasoning patterns contained in a sample, independent of the sample’s surface length.
Figure 4: Experimental results of fine-tuning models in different intervals. "Direct-All" denotes a model fine-tuned on all training examples. "Range(x, y)" denotes a model fine-tuned on examples whose SLKD scores fall between $x$ and $y$. Left: Test dataset experiment results. Right: Balanced testing set experiment results. The horizontal axis represents score intervals, and the vertical axis represents error rates (lower error indicates better performance).
Figure 5: Experimental results of fine-tuning models using different methods. "Ours-All": model fine-tuned on the full training set using re-cognizing optimization. "Ours-Range(x, y)": model fine-tuned with re-cognizing optimization on examples whose SLKD scores fall between $x$ and $y$. Left: Test dataset results. Right: Balanced testset results. The horizontal axis represents score intervals and the vertical axis represents error rates (lower error indicates better performance).
4 Empirical Analysis for SLKD
In order to fully investigate the effect of SLKD, we conduct experiments in existing training data (experimental setting is in Appendix C). We not only focus on the SLKD score distribution of the existing data but also use data with different interval distributions as training sets to separately validate the effectiveness of the data within these intervals.
4.1 SLKD Score Distribution
In Figure 3(a), when the training set is binned into 20 score intervals, the sample counts form a bell-shaped curve with a weighted mean of $\mu=0.526$ and standard deviation $\sigma=0.204$. This approximately Gaussian frequency distribution demonstrates that our SLKD scores effectively separate examples by their underlying reasoning potential. Figure 3(a) also plots model error rate against these intervals: error climbs from 8.5% at a score of 0.1 to 59.3% at 0.6, demonstrating a clear positive correlation between SLKD and failure rate. Beyond a score of 0.6, error rate plateaus at approximately 61% ± 3.6%. This saturation likely arises because (1) very high SLKD items exceed the model’s current reasoning capacity, and (2) the top bins contain few, uniformly hard cases, so failures become uniformly pervasive. Unlike surface metrics (e.g., sentence length or vocabulary complexity), our score distribution captures deeper logical structure.
4.2 SLKD Score Distribution Balancing
The original testset contains 4,743 examples and its SLKD score distribution (Figure 3(b)) reveals that 74.3% of samples fall into the mid-range interval (0.2, 0.7), while only 17.2% lie in the high SLKD interval (>0.7). Such skew can introduce two evaluation biases: (1) Apparent overfitting in the mid-range: strong results in the overrepresented middle interval may conceal weaknesses on higher SLKD examples; (2) Insufficient statistical power: the small number of high SLKD samples ($n=818$) yields wide confidence intervals for any observed gains. To correct this, we build a balanced benchmark (Figure 3(c)) by drawing 80 examples from each interval. This uniform sampling ensures equal representation across the SLKD spectrum, eliminating bias and enabling reliable comparisons.
4.3 Data Potential Analysis by SLKD
Building on our validation of the SLKD score and the creation of a balanced test benchmark, we next investigated how filtering training examples by their SLKD scores affects learning. In Figure 4, we systematically evaluate model performance when trained on subsets defined by SLKD intervals, using both the original and balanced test sets. These experiments reveal three critical patterns, which can be summarized as the following observations: